OpenStudy (anonymous):

A machinist creates a washer by drilling a hole through the center of a circular piece of metal. If the piece of metal has a radius of x + 10 and the hole has a radius of x + 6, what is the area of the washer?

OpenStudy (anonymous):

@hal_stirrup im here now ;)

OpenStudy (anonymous):

hi

OpenStudy (anonymous):

my last one

OpenStudy (anonymous):

ok what is the area of hole?? right.

OpenStudy (anonymous):

6

OpenStudy (anonymous):

A(washer) - A(hole) = π (x+10)^2 - π (x+6)^2 = π (8x+64)

OpenStudy (anonymous):

@dan18 would you rathe help daboss33 understand how yo think about the problem instead of solving it.!!

OpenStudy (anonymous):

@hal_stirrup are you mad? sorry I thought it would be much more easier to understand since the solution is given already, the solution I've given speak itself. A(washer) - A(hole) = πR^2 - πr^2 = π(x+10)^2 - π(x+6)^2 =π(x^2+20x+100) - π(x^2+12x+36) = π(8x+64) sq. unit